OpenStudy (anonymous):

My Calc is a little rusty: Why is lim k-> infinity (k+1)/k =1? and lim k-> infinity (2^(k+1))/2k = 2?

OpenStudy (anonymous):

same as a horizontal asymptote the degrees are the same, the limit is the ratio of the leading coefficients

OpenStudy (kkutie7):

trying plugging in bigger and bigger numbers into your function for k.... take the derivative of the bottom and top (1/1)=1

OpenStudy (anonymous):

for the first one anyways

zepdrix (zepdrix):

For the second one, did you mean to type: lim k-> infinity (2^(k+1))/2^k = 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Imagine k very large, like 10^6, and take the ratio, which becomes almost exactly 1 for the second try logs (k+1) log 2 -log 2 - log k and let k get very large, leaving k log 2 taking anti logs we have 2^k , which may be what you were looking for.

zepdrix (zepdrix):

For the second one I think maybe you're just forgetting your rules of exponents yes?\[\Large\rm 2^{k+1}=2^k\cdot 2^1\]So that second limit becomes:\[\Large\rm \frac{2\cdot 2^k}{2^k}\]And then before taking your limit, you have a nice cancellation.

OpenStudy (anonymous):

oh yes I did. Thanks so much!

OpenStudy (anonymous):

The denominator was 2^k ??? Oh, well.

OpenStudy (anonymous):

yeah sorry about that typo